OpenStudy (yrivers36):

Please help me with this system. I'm not getting the y or z. -x+2y+z=2 3x+y+z=12 x-y-2z=-1

OpenStudy (anonymous):

(3,2,1)

OpenStudy (yrivers36):

Can u explain it to me please?

OpenStudy (anonymous):

are you supposed to use elimation?

OpenStudy (yrivers36):

Either elimination or substitution

OpenStudy (anonymous):

ok we can eleminate z first

OpenStudy (anonymous):

satellite, i wasnt done, i still needed helpp :/

OpenStudy (anonymous):

\[-x+2y+z=2\] \[ 3x+y+z=12\]

OpenStudy (anonymous):

subtract and get \[-4x+y=-10\] as one equation with no z in it

OpenStudy (anonymous):

now \[3x+y+z=12\] \[ x-y-2z=-1\] multiply first equation by 2 get \[6x+2y+2z=24\] \[x-y-2z=-1\] add and get \[7x+y=23\]

OpenStudy (anonymous):

now we have \[-4x+y=-10\] and \[7x+y=23\] so solve

OpenStudy (anonymous):

subtact second one from first and get \[-11x=-33\] making \[x=3\] now substitute back to find y, and then z

OpenStudy (yrivers36):

Where did the y go in the -4x+y=-10 and 7x+y=23

OpenStudy (anonymous):

i subtracted the second equation from the first

OpenStudy (yrivers36):

oh you multiplied by -1

OpenStudy (anonymous):

or if you prefer, multiply the second by -1 and then add,

OpenStudy (anonymous):

yes exactly it is the same thing

OpenStudy (yrivers36):

ok I got it finally thank you so much for your help

OpenStudy (anonymous):

\[-4x+y=-10\] \[-7x-y=-23\] \[-11x=-33\] etc yw